OpenStudy (anonymous):

You make an initial investment of $5,000 in an account that pays 6.75% interest compounded semi-annually. How long will it take for the account to have $7,500 in it?

OpenStudy (jdoe0001):

http://openstudy.com/updates/538510e6e4b07cf9dcd34620 <--- same way, same formula

OpenStudy (anonymous):

but how do you do that?

OpenStudy (jdoe0001):

same way as show on the other one you'd use the same "compound interest" formula

OpenStudy (jdoe0001):

just keep in mind the cycle period -> compounded semi-annually.

OpenStudy (anonymous):

A + P (1 + r/n)^rt ?

OpenStudy (jdoe0001):

hmmm actually.... I see the issue... you have to solve for "t" or years.... ok have you covered logarithms yet?

OpenStudy (anonymous):

kind of, I'm still a little confused with it

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt}\implies \cfrac{A}{P}=\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ \textit{log cancellation rule of }log_{\color{red}{ a}}({\color{red}{ a}}^x)=x\qquad thus \\ \quad \\ \large log_{\left(1+\frac{r}{n}\right)}\frac{A}{P}= log_{\left({\color{brown}{ 1+\frac{r}{n}}}\right)}\left[{\color{brown}{ 1+\frac{r}{n}}}\right]^{nt}\implies log_{\left(1+\frac{r}{n}\right)}\frac{A}{P}=nt \\ \quad \\ \large \cfrac{log_{\left(1+\frac{r}{n}\right)}\frac{A}{P}}{n}=t\)

OpenStudy (jdoe0001):

well \(\bf A=P\left(1+\frac{r}{n}\right)^{nt}\implies \cfrac{A}{P}=\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ \textit{log cancellation rule of }log_{\color{red}{ a}}({\color{red}{ a}}^x)=x\qquad thus \\ \quad \\ \large log_{\left(1+\frac{r}{n}\right)}\left[\frac{A}{P}\right]= log_{\left({\color{brown}{ 1+\frac{r}{n}}}\right)}\left[{\color{brown}{ 1+\frac{r}{n}}}\right]^{nt}\implies log_{\left(1+\frac{r}{n}\right)}\left[\frac{A}{P}\right]=nt \\ \quad \\ \large \cfrac{log_{\left(1+\frac{r}{n}\right)}\left[\frac{A}{P}\right]}{n}=t\)